# OCRBench v2: An Improved Benchmark for Evaluating Large Multimodal Models on Visual Text Localization and Reasoning

### v1

@@ 5 Conclusion
In this work, we introduce OCRBench v2, an improved benchmark for assessing the OCR capabilities of LMMs. By studying $38$ representative LMMs across $23$ OCR tasks, we not only reveal some defects of the current LMMs but also explore factors affecting the OCR performance of LMMs. We hope OCRBench v2 could facilitate future research on improving the OCR capabilities of LMMs.
Limitations. One challenge we encountered is that LMMs sometimes produce responses that deviate from the given instructions, making it difficult to extract the desired answers. In future work, we plan to develop a more objective assessment framework to address this issue.
Acknowledgements
We thank Max Tyrrell for providing valuable receipt images, which have significantly contributed to the development of OCRBench v2.
Appendix
This appendix mainly contains the following contents:
  • Section A: Comparison experiments between LMMs and some text-centric expert models.
  • Section B: Data source of OCRBench v2.
  • Section C: Task definition of OCRBench v2.
  • Section D: More statistics of OCRBench v2.
  • Section E: Evaluation metrics of OCRBench v2.
  • Section F: Visualization samples of task examples.
  • Section G: Visualization samples of failure cases.
Appendix A Comparison with LMMs and Expert Models
Comparison with text recognizers. We compare LMMs with several representative scene text recognizers, including CRNN [87], ABINet [23], ASTER [89], MASTER [63], and SVTR [22], on the text recognition task. The weights of these models are loaded from mmocr [^3] https://github.com/open-mmlab/mmocr. The results are shown in Table 6, where we selected the top three LMMs with the best performance, including Qwen2VL-8B [105], GPT4V [78] and Step-1V [94]. The results demonstrate that LMMs exhibit remarkable text recognition capabilities, validating our motivation to evaluate LMMs on more challenging OCR-related tasks.
@@ Appendix A Comparison with LMMs and Expert Models
Comparison with text spotters. We also compare LMMs with ABCNet series [56, 57] and TESTR [130] on the text spotting task. The weights of the ABCNet series are loaded from mmocr and TESTR is loaded from the official checkpoint [^4] https://github.com/mlpc-ucsd/TESTR fine-tuned with TotalText [15]. The results are shown in Table 7. We observe that although LMMs demonstrate promising capabilities in text recognition, there remains significant potential for improvement in the text spotting task.
Comparison with GOT. We noticed a recent work, GOT [109], that can parse the textual elements within images. We conduct comparison experiments between GOT and some representative LMMs, and the results are shown in Table 8. We observe that LMM shows advantages in general text recognition, while GOT demonstrates better performance in document parsing task.
@@
Appendix B Data Collection
Text Recognition. The data for text recognition task are sampled from ICDAR2013 [38], SVT [91], IIIT5K [75], ICDAR2015 [39], SCUT-CTW1500 [55], COCO-Text [102], CUTE80 [84], TotalText, SVTP [83], WordArt [112], NonSemanticText [58], IAM [68], ORAND-CAR-2014 [19], HOST [108] and WOST [108]. Meanwhile, CAPTCHA(Completely Automated Public Turing Test to Tell Humans Apart) images used in text recognition are sourced from a CAPTCHA dataset [^5] https://aistudio.baidu.com/datasetdetail/159309 and a number CAPTCHA dataset [^6] https://www.heywhale.com/mw/dataset/5e5e56b6b8dfce002d7ee42c/file. And dot matrix images in the text recognition task are obtained from the web page.
Fine-grained Text Recognition. In the fine-grained text recognition task, the images are sampled from the testset of Fox [50], Totaltext, COCO-Text, CTW1500 [127] and ICDAR2015. For the Fox dataset, we used the original annotations provided. For the other four datasets, we performed manual annotation.
@@ Appendix B Data Collection
VQA with Position. The images for VQA with position task are sampled from testset of TextVQA [93] and RICO [18]. QA pairs and bounding boxes annotations are based on their official OCR annotations.
Text Spotting. The data sources for text spotting task include Totaltext, COCO-Text, CTW1500, and ICDAR2015.
Key Information Extraction. The data sources for key information extraction task include FUNSD [35], SROIE [34], POIE [43], M6Doc, XFUND [113], ICDAR2023-SVRD [123] and a private receipt photograph dataset.
@@
Handwritten Content Extraction. This task’s data is our private data, which contains real exam paper data with student information removed and manually annotated question-answer pairs.
@@
Document Parsing. The data sources for document parsing task come from DoTA [49], DocVQA [72], M6Doc and CDLA.
Formula Recognition. The data sources for the formula Recognition task including HME100K [125], IM2LATEX-100K [104], M2E [115], MathWriting [27], MLHME-38K [^9] https://ai.100tal.com/icdar, CASIA-CSDB [20] and some private data.
Math QA. The data sources for the math QA task including MathMatics [86], MathVerse [129], MathVision [104] and MathVista [65].
@@
Cognition VQA. The data sources for the cognition VQA task include EST-VQA [107], OCRVQA [76], ST-VQA [7], TEXTVQA, DIR300 [24], ChartQA [69], DVQA [37], PlotQA [74], InfoVQA [70], WTW, PubTabNet [136], WTQ [82], CORD [81], LLaVAR [131], WebSRC [11], DocVQA, M6Doc, XFUND, Publaynet [137], RVL-CDIP [29], ScreenQA [4], SlideVQA [97], a movie poster collection dataset [^10] https://www.kaggle.com/datasets/neha1703/movie-genre-from-its-poster, a website screenshot collection dataset [^11] https://huggingface.co/datasets/Zexanima/website_screenshots_image_dataset/tree/main and a private receipt photograph dataset.
@@
Document Classification. The images for the document classification task are mainly collected from RVL-CDIP.
@@
Text Translation. The datasets collected for text translation task includes memes [^12] https://www.kaggle.com/datasets/dvishal485/meme-challenge?resource=download, MSRA-TD500 [116], MTWI2018 [30], M6Doc, ICDAR2023-SVRD, EST-VQA, RCTW17 [90], DAST1500 [98], XFUND, ArT2019 [16], ChartQA, CDLA, ICDAR2015, SlideVQA, Fintabnet [135], ScienceQA, InfoVQA, COMICS-Dialogue [^13] https://huggingface.co/datasets/lmms-lab/M4-Instruct-Data and ExpressExpense SRD [^14] https://expressexpense.com/blog/free-receipt-images-ocr-machine-learning-dataset/.
Appendix C Task Definition
In this section, we introduce the definition of each task, and the visualizations for each task can be found in Section F.
@@ Appendix C Task Definition
Table 9: The number of images included in each scene category.
Figure 8: The quantity distribution of English tasks.
Figure 9: The quantity distribution of Chinese tasks.
Figure 10: The OCR lines distribution of English tasks.
Figure 11: The OCR lines distribution of Chinese tasks.
@@
Document Parsing. In the document parsing task, both text and the complex elements, including chart, table, and formula, are required to parse.
@@
Science QA. In the Science QA task, LMMs are required to respond to the scientific problem. We use PaddleOCR [^15] https://github.com/PaddlePaddle/PaddleOCR to extract text from the collected images and filter out those with fewer than four OCR results. Additionally, when extra subject-related knowledge is provided by the source, we incorporate them by rendering them into the images.
@@
Appendix D More Statistics of OCRBench v2
D.1 Scene Coverage
Our dataset can be divided into 31 classic scenes according to the scene of the image. The specific scenes and the corresponding number of pictures are shown in Table 9.
D.2 Statistics of each task
The distribution of 23 tasks in OCRBench v2 is displayed in Figure 8 and Figure 9. Additionally, we calculate and present the average number of OCR text lines per task in Figure 10 and Figure 11. As illustrated in these figures, the task distribution is well-balanced, with each task containing adequate textual information for analysis.
Appendix E Evaluation Metrics
Parsing Type. We use Tree-Edit-Distance-based Similarity (TEDS) [138] to evaluate parsing tasks, which require LMMs to transform the images to structured formats. Tree Edit Distance (TED) refers to the minimum number of edits to transform one tree into another. TEDS is based on TED to calculate the similarity of two trees. Assuming $T_{1}$ and $T_{2}$ are two different trees, $TED(T_{1},T_{2})$ refer to their TED, the TEDS is defined as:
@@ Appendix E Evaluation Metrics
where $|T_{1}|$ and $|T_{2}|$ is the number of nodes of trees, $TED(T_{1},T_{2})$ can be calculated by dynamic programming algorithm. If $T_{1}$ and $T_{2}$ are identical, then their TEDS equals to $1$. As the structural difference between two trees increases, their TED value becomes larger, resulting in the TEDS approaching 00.
@@
Long Reading Type. To evaluate LMMs’ ability to recognize text across entire paragraphs or pages, BLEU [80], METEOR [6], F1 score and normalized edit distance are employed. And the final score is the average value of these metrics.
BLEU evaluates prediction quality by comparing n-gram match rates between prediction and ground truth sequences. For each n-gram type, precision is calculated as the ratio of matching n-grams to total prediction n-grams. The final BLEU score is the geometric mean of these precision values multiplied by a penalty $BP$, which is defined as:
@@
METEOR employs a semantic-aware matching strategy with four levels. 1) Exact Match: words in prediction that are identical to ground truth. 2) Stem match: matching words that have the same word stem. 3) Synonym Match: matching words based on synonymous relationships. 4) Paraphrase Match: Matching similar phrases at the phrase level. These matches are combined to calculate precision and recall, from which a weighted harmonic mean F1 score is derived as:

### v2

@@ 6 Conclusion
In this work, we introduce OCRBench v2, a comprehensive benchmark designed to evaluate the OCR capabilities of LMMs. Covering $23$ tasks across $31$ diverse scenarios, our benchmark systematically assesses eight core capabilities that are essential for text-oriented visual understanding tasks. It includes $10,000$ high-quality QA pairs and six rigorous evaluation metrics. In addition, we curate a private test set of $1,500$ manually labeled images to ensure robust generalization evaluation. Leveraging this benchmark, we conduct extensive experiments on representative LMMs. Through in-depth analysis of experimental results, we identify critical limitations of current models and uncover key factors that affect their OCR performance. We hope OCRBench v2 could aid future research on enhancing LMMs’ text understanding ability.
Appendix A Technical Appendices and Supplementary Material
This supplementary material contains the following content:
  • Sec. A.1: Comparison experiments between LMMs and some text-centric expert models.
  • Sec. A.2: Data collection.
  • Sec. A.3: Task definitions.
  • Sec. A.4: Additional statistics of OCRBench v2.
  • Sec. A.5: Evaluation metrics.
  • Sec. A.6: Experimental setting for the evaluation process.
  • Sec. A.7: Compute resources for the evaluation process.
  • Sec. A.8: Evaluation results for LMMs on OCRBench v2.
  • Sec. A.9: Potential factors affecting OCR capabilities
  • Sec. A.10: Visualization samples for task examples.
  • Sec. A.11: Visualization samples for failure cases.
  • Sec. A.12: Discussion of broader impacts.
  • Sec. A.13: Discussion of limitations.
A.1 Comparison with LMMs and Text-centric Expert Models
Comparison with text recognizers. We compare LMMs with several representative scene text recognizers, including CRNN [67], ABINet [68], ASTER [69], MASTER [70], and SVTR [71], on the text recognition task. The weights of these models are loaded from mmocr [^1] https://github.com/open-mmlab/mmocr. The results are shown in Tab. 6, where we selected $5$ representative LMMs, including Qwen2.5VL-7B [66], InternVL3-14B [23], GPT4o [1], Gemini1.5-Pro [60], and Step-1V [62]. The results demonstrate that LMMs exhibit remarkable text recognition capabilities, validating our motivation to evaluate LMMs on more challenging OCR-related tasks.
@@ A.1 Comparison with LMMs and Text-centric Expert Models
Comparison with text spotters. We also compare LMMs with ABCNet series [72, 73] and TESTR [74] on the text spotting task. The ABCNet series utilize the official weights [^2] https://github.com/aim-uofa/AdelaiDet, and TESTR is also initialized with its publicly released checkpoint [^3] https://github.com/mlpc-ucsd/TESTR. These models were fine-tuned with TotalText [75]. The results are shown in Tab. 7. Although LMMs demonstrate promising capabilities in text recognition, there remains notable potential for improvement in the text spotting task.
@@
Comparison with GOT. We notice a recent work, GOT [76], that can parse the textual elements within images. We conduct comparison experiments between GOT and some representative LMMs, and the results are shown in Tab. 8. We observe that LMMs show advantages in general text recognition, while GOT demonstrates better performance in the document parsing task.
@@
A.2 Data Collection
Text Recognition. The data for text recognition task are sampled from ICDAR2013 [77], SVT [78], IIIT5K [79], ICDAR2015 [80], SCUT-CTW1500 [81], COCO-Text [82], CUTE80 [83], TotalText, SVTP [84], WordArt [85], NonSemanticText [14], IAM [86], ORAND-CAR-2014 [87], HOST [88], and WOST [88]. Meanwhile, CAPTCHA (Completely Automated Public Turing Test to Tell Humans Apart) images are sourced from a CAPTCHA dataset [^4] https://aistudio.baidu.com/datasetdetail/159309 and a number CAPTCHA dataset [^5] https://www.heywhale.com/mw/dataset/5e5e56b6b8dfce002d7ee42c/file. Additionally, dot matrix images in the text recognition task are manually collected from the web page.
Fine-grained Text Recognition. In the fine-grained text recognition task, images are sampled from the test sets of Fox [17], Totaltext, COCO-Text, CTW1500 [89], and ICDAR2015. We use the original annotations for Fox, while the other datasets are manually re-annotated.
@@ A.2 Data Collection
VQA with Position. The images used for VQA with position task are sampled from the test sets of TextVQA [10] and RICO [96], with QA pairs and bounding box annotations derived from their original datasets.
Text Spotting. The data sources for the text spotting task include Totaltext, COCO-Text, CTW1500, and ICDAR2015.
Key Information Extraction. The data sources for key information extraction task include FUNSD [97], SROIE [98], POIE [99], M6Doc, XFUND [100], ICDAR2023-SVRD [101], and a private dataset of photographed receipts.
@@
Handwritten Content Extraction. This task’s data is our private data, which contains real exam paper data with student information removed and manually annotated QA pairs.
@@
Document Parsing. The data sources for document parsing task come from DoTA [104], DocVQA [105], M6Doc, and CDLA.
Formula Recognition. The data sources for the formula Recognition task includes HME100K [106], IM2LATEX-100K [107], M2E [108], MathWriting [109], MLHME-38K [^8] https://ai.100tal.com/icdar, CASIA-CSDB [110], and some private data.
Math QA. The data sources for the math QA task includes MathMatics [111], MathVerse [112], MathVision [107], and MathVista [113].
@@
Cognition VQA. The data sources for the cognition VQA task include EST-VQA [12], OCRVQA [114], ST-VQA [11], TEXTVQA, DIR300 [115], ChartQA [41], DVQA [116], PlotQA [117], InfoVQA [118], WTW, PubTabNet [119], WTQ [120], CORD [121], LLaVAR [30], WebSRC [122], DocVQA, M6Doc, XFUND, Publaynet [123], RVL-CDIP [124], ScreenQA [125], SlideVQA [126], a movie poster collection dataset [^9] https://www.kaggle.com/datasets/neha1703/movie-genre-from-its-poster, a website screenshot collection dataset [^10] https://huggingface.co/datasets/Zexanima/website_screenshots_image_dataset/tree/main, and a private receipt photograph dataset.
@@
Document Classification. The images for the document classification task are collected from RVL-CDIP.
@@
Text Translation. The datasets collected for text translation task includes memes [^11] https://www.kaggle.com/datasets/dvishal485/meme-challenge?resource=download, MSRA-TD500 [134], MTWI2018 [135], M6Doc, ICDAR2023-SVRD, EST-VQA, RCTW17 [136], DAST1500 [137], XFUND, ArT2019 [138], ChartQA, CDLA, ICDAR2015, SlideVQA, Fintabnet [139], ScienceQA, InfoVQA, COMICS-Dialogue [^12] https://huggingface.co/datasets/lmms-lab/M4-Instruct-Data, and ExpressExpense SRD [^13] https://expressexpense.com/blog/free-receipt-images-ocr-machine-learning-dataset/.
A.3 Task Definitions
In this section, we introduce the definition of each task, and the visualizations for each task can be found in Sec. A.10.
@@ A.3 Task Definitions
Table 9: The number of images included in each scene category in public data.
Figure 6: The quantity distribution of English tasks of public data.
Figure 7: The quantity distribution of Chinese tasks of public data.
Figure 8: The OCR lines distribution of English tasks of public data.
Figure 9: The OCR lines distribution of Chinese tasks of public data.
@@
Document Parsing. In the document parsing task, both text and the complex elements, including charts, tables, and formulas, are required to be parsed.
@@
Science QA. In the Science QA task, LMMs are required to respond to the scientific problem. We use PaddleOCR [^14] https://github.com/PaddlePaddle/PaddleOCR to extract text from the collected images and filter out those with fewer than four OCR results. Additionally, when extra subject-related knowledge is provided by the source, we incorporate it by rendering it into the images.
@@
A.4 Additional Statistics of OCRBench v2
Scene Coverage. Our dataset can be divided into $31$ classic scenes according to the scene of the image. The specific scenes and the corresponding number of pictures are shown in Tab. 9.
Statistics of each task. Fig. 5 shows an overview of each task in OCRBench v2.The distribution of $23$ tasks in OCRBench v2 is displayed in Fig. 6 and Fig. 7. Additionally, we calculate and present the average number of OCR text lines per task in Fig. 8 and Fig. 9. As illustrated in these figures, the task distribution is well-balanced, with each task containing adequate textual information for analysis.
A.5 Evaluation Metrics
Parsing Type. We use Tree-Edit-Distance-based Similarity (TEDS) [63] to evaluate parsing tasks, which require LMMs to transform the images to structured formats. Tree Edit Distance (TED) refers to the minimum number of edits to transform one tree into another. TEDS is based on TED to calculate the similarity of two trees. Assuming $T_{1}$ and $T_{2}$ are two different trees, $TED(T_{1},T_{2})$ refers to their TED, and the TEDS is defined as:
@@ A.5 Evaluation Metrics
where $|T_{1}|$ and $|T_{2}|$ is the number of nodes of trees, $TED(T_{1},T_{2})$ can be calculated by dynamic programming algorithm. If $T_{1}$ and $T_{2}$ are identical, then their TEDS equals $1$. As the structural difference between two trees increases, their TED value becomes larger, resulting in the TEDS approaching 00.
@@
Long Reading Type. To evaluate LMMs’ ability to recognize text across entire paragraphs or pages, BLEU [64], METEOR [65], F1 score, and normalized edit distance are employed. And the final score is the average value of these metrics.
BLEU evaluates prediction quality by comparing n-gram match rates between the prediction and ground truth sequences. For each n-gram type, precision is calculated as the ratio of matching n-grams to total predicted n-grams. The final BLEU score is the geometric mean of these precision values multiplied by a penalty $BP$, which is defined as:
@@
METEOR employs a semantic-aware matching strategy with four levels. 1) Exact Match: words in the prediction that are identical to the ground truth. 2) Stem match: matching words that have the same word stem. 3) Synonym Match: matching words based on synonymous relationships. 4) Paraphrase Match: Matching similar phrases at the phrase level. These matches are combined to calculate precision and recall, from which a weighted harmonic mean F1 score is derived as: